# Adaptation to extreme stress under the growth-survival fitness trade-off

### v1

@@ 3 Long Term Growth Rate and Comparing Fitness
where $T_{s}$ is the duration of stress exposure and $d$ is the dilution rate. Note that we do not explicitly account for transition to quiescence here for simplicity of calculation, which means that $s(\phi)$ does not act upon a smaller quiescent population, but rather on the entire population. The survival process in this calculation can be thought of as an effective process that combines quiescence and survival from our simulations. After $c$ cycles, then, we have a population distribution
@@
Figure 4 shows a comparison of the equilibrium phenotype values from the simulation with the optimal phenotype that maximized the long term growth rate in equation 8. As was seen in the last section, the equilbrium phenotype increases with both an increase in $g_{0}$ and $\tau_{l}$. Note that we made the assumption above that the total time of growth, $T$, remains constant across cycles. This assumption does not hold in the simulations as the time of growth decreases with increasing population size as the evolution progresses forward. Growth time in our simulations drops from a relatively high value in the beginning to a lower equilibrium value. For the growth time dynamics, see Supplementary Information. We use the mean of the growth time from simulations as $T$ in 8 to calculate the optimum phenotype. Since we are comparing a finite resource growth process (in the simulations) to a finite time growth process (in the analytic calculation), some difference in the two predictions is seen. However, the analytical prediction agrees with the values from the simulation to a large degree.
@@
Note that when the long term growth rate is negative, the strain is expected to go extinct. As can be seen, for most parameter pairs, the performance of the population remains in the upper right quadrant under growth cycles (circular points), i.e. populations of both $\phi_{G0}$ and $\phi_{S0}$ grow well under only growth cycles. When each circular point is compared with its counterpart under growth-stress cycles (diamond points), you see a shift towards lower $\Lambda$ for both phenotypes.
@@
Some populations, with larger $g_{0}$ values, remain in the upper right quadrant, implying both $\phi_{G0}$ or $\phi_{S0}$ are able to grow under both under growth-stress, only with reduced rates (with $\Lambda_{i}(\phi_{S0})$ typically decreasing less than $\Lambda_{i}(\phi_{G0})$. Thus, two populations of $\phi_{G0}$ and $\phi_{S0}$ are generalists with different emphasis on growth and survival. This implies that in cases where the population variance is large enough, coexistence may be possible.
Some populations under growth-stress cycles shift to the lower left quadrant, implying neither $\phi_{G0}$ or $\phi_{S0}$ are able to grow under stress. The counterparts for these populations under solely growth cycles are in the first quadrant, implying that both $\phi_{G0}$ or $\phi_{S0}$ are able to grow, and even possibly coexist. This occurs at lower values of $g_{0}$, where the low fecundity does not allow the population to recover its numbers after the significant decrease in its population size upon exposure to stress. Increasing both $g_{0}$ and $T$ can help the population recover and reduce this effect.
@@
A few populations under growth-stress fall in the second quadrant, showing survival of $\phi_{S0}$ under stress conditions and extinction of $\phi_{G0}$. In other words, a strain evolved under stressful conditions survives under growth-stress cycles, while one evolved under growth cycles does not. However, the counterparts of these populations under solely growth cycles exist in the upper right quadrant. This shows that even though a population of $\phi_{S0}$ evolves to exist under stress conditions when $\phi_{G0}$ cannot, under solely growth cycles the two populations may coexist. Hence, despite the growth-survival trade-off, populations evolved under growth-stress may not show significant trade-offs with their fitness under only growth cycles.
4 Discussion
@@
Growth-survival trade-offs, observed across microbial species, help explain key ecological phenomena like co-existence, population heterogeneity, and the preference for resource-scarce environments [36, 37]. We present a model of adaptation to extreme stress under such a trade-off. To our knowledge, this is a novel effort to model the role of quiescence in evolutionary adaptation to extreme stress. This approach is supported by evidence that stress response in yeast may follow a general strategy rather than evolving to specific stress conditions [1, 20, 38]. Studies in bacteria suggest that exposure to lower concentrations of antibiotics may prepare cells to take on the challenge of lethal concentrations [39, 40]. Further, the existence of the growth-lag trade-off suggests that starvation in stationary phase may be perceived a mild stress [41]. These observations support our assumption that quiescence, a response to the relatively mild stress of starvation, may provide a helpful step towards resistance to harsher stress.
@@
Previous work has demonstrated that yeast populations increase their survival from $2\%$ to $70\%$ over the course of $25$ generations [9]. These evolved strains show key characteristics of quiescent cells including smaller, denser cells, increased cell wall thickness. Our model captures this evolutionary trajectory qualitatively at the phenotypic level. It also provides a framework to explore its dependence on the frequency and strength of environmental change, as well as on features of the population such as its maximum growth rate and the nature of the variation of lag time with the population phenotype.
@@
We model mutational effects drawn from a long-tailed Pearson distribution. The model is agnostic to the source of such mutations, which could be genetic or epigenetic in nature. Further, the framework does not speculate on the mechanism through which the memory of stress, or of intracellular trehalose produced in stationary phase can be inherited across generations. Interestingly, the whole genome sequencing of the evolved yeast strains in [9] does not show any mutations conserved across replicate lines. This is also supported by previous work [47]. Yet, a convergent phenotype emerges within the short period of $25$ generations, suggesting an epigenetic stress response. This response seems to still be inheritable, as evidenced by the evolved phenotype persisting even in the absence of continued stress. It is possible that if quiescence, a generalized stress response mechanism, facilitates the adaptation to freeze thaw, that the adaptation mechanism is one that involves genetic assimilation [48, 49].
We assume a stochastic switch to quiescence, with the probability of quiescence increasing with increasing intracellular trehalose. In this way our system is similar to stochastic switching to a persistent state in bacterial populations, although for us probability of switching is biased by trehalose concentration in stationary phase [50, 51, 52]. However, our model is distinct from most models of bacterial persistors since we postulate that our phenotype is inheritable and over time lends higher survival fractions. There is, however, some support for the idea that that persistence frequency in bacteria is also an evolvable trait that adapts to both the duration and frequency of the antibiotic application[53, 54]. We assume that the subset of quiescent cells is then subject to probabilistic survival upon exposure to stress. This is a reasonable assumption given previous evidence that quiescence leads to higher stress tolerance [55]. Yet this assumption remains experimentally unconfirmed, and it is not certain whether the same quiescent subset of the population in stationary phase is selected from by subsequent stress.
@@
We calculate the long term growth rate of a population adapting to growth-stress cycles and use it to calculate the optimum phenotype for different evolution programs. Our analysis reveals that the duration and nature of the growth process affects directly the optimum phenotype, final survival, as well as the long term growth rate. We are also able to compare the fitnesses of populations that are optimum under growth cycles and under growth-stress cycles. We calculate the long term growth rate of these ’growth-type’ and ’stress-type’ populations both with and without stress. We find that depending on the evolution program, the populations can show the possibility of coexistence, or show a trade-off in their overall fitness. The populations show the possibility of coexistence both with and without stress at large fecundity, $g_{0}$, indicating that their overall fitness may not show a strong trade-off if they are able to recover their populations in the growth phase after exposure to stress, where the population size decreases drastically.
For a smaller range of lower values of $g_{0}$ and time of growth, $T$, we see that ’stress-type’ populations may survive both types of environments, while ’growth-type’ populations do not survive cycles with stress exposure. Thus, for some evolution paradigms, populations that do not show a significant decrease in their fitness in growth-only cycles while still increasing their fitness in cycles with stress exposure exist. This is because the overall fitness and long term growth rate depend not only on the relative strengths of the growth and survival functions but also on other characteristics of growth such as time, $T$ and maximum growth rate $g_{0}$. Several experimental studies have shown that fitness trade-offs are not ubiquitous in microbial systems [4, 56]. Our results demonstrate that the absence of more obvious fitness trade-offs does not preclude the existence of an underlying growth-survival trade-off.
@@
Our model is also reminiscent of recent work done on the population dynamics of bacteria in stochastic feast-famine cycles [58, 59]. Previous work which models growth and death processes both with growth-yield and growth-death trade offs shows that the ratio of the growth rate and death rate is an effective fitness of the population [58]. This is similar to our expression for the long term growth rate which combines survival probability with the growth rate. Further, they also find the duration of the feast and famine periods to be important to the eventual fate of the population. However, here we also take into account the added complexity of lag time along with growth and survival, all determined by a single phenotype. Other work has examined the optimal persistence strategy for bacteria in feast-famine cycles with the stochastic application of antibiotics [60] to find that ’triggered’ dormancy, such as what we have included in this model with quiescence, may be the optimal strategy when the stress of the antibiotic treatment is very severe.

### v2

@@ III Long Term Growth Rate and Comparing Fitness
where $T_{s}$ is the duration of stress exposure and $d$ is the dilution rate. Note that we do not explicitly account for transition to quiescence here for simplicity of calculation, which means that $s(\phi)$ does not act upon a smaller quiescent population, but rather on the entire population. The survival process in this calculation can be thought of as an effective process that combines quiescence and survival from our simulations. After $c$ cycles we have a population distribution
@@
Figure 5 shows a comparison of the equilibrium phenotype values from the simulation with the optimal phenotype that maximized the long term growth rate in equation 8. As was seen in the last section, the equilibrium phenotype increases with both an increase in $g_{0}$ and $\tau_{l}$. Note that we made the assumption above that the total time of growth, $T$, remains constant across cycles. This assumption does not hold in the simulations as the time of growth decreases with increasing population size as the evolution progresses forward. Growth time in our simulations drops from a relatively high value in the beginning to a lower equilibrium value. For the growth time dynamics, see Supplementary Information. We use the mean of the growth time from simulations as $T$ in 8 to calculate the optimum phenotype. Since we are comparing a finite resource growth process (in the simulations) to a finite time growth process (in the analytic calculation), some difference in the two predictions is seen. However, the analytical prediction agrees with the values from the simulation to a large degree.
@@
When the long term growth rate is negative, the strain is expected to go extinct quickly. For most parameter pairs, the performance of the population remains in the upper right quadrant under growth cycles (circular points), i.e. populations of both $\phi_{G0}$ and $\phi_{S0}$ grow well under only growth cycles. This is striking since it shows that although $\phi_{S0}$ is evolved to be optimum under growth-stress cycles, it can perform well under solely growth cycles. Further, $\phi_{G0}$ and $\phi_{S0}$ show comparable long term growth rates, which points to the possibility of coexistence. Thus, despite the strong growth-survival trade-off, the overall fitness does not show a strictly inhibiting trade-off for $\phi_{S0}$ under growth conditions. When each circular point is compared to its counterpart in growth-stress cycles (diamond points), a shift is observed towards lower $\Lambda$ for both phenotypes.
@@
Some populations, with larger $g_{0}$ values, remain in the upper right quadrant, implying both $\phi_{G0}$ or $\phi_{S0}$ are also able to grow with positive long term growth rates under under growth-stress cycles, only with reduced rates (with $\Lambda_{i}(\phi_{S0})$ typically decreasing less than $\Lambda_{i}(\phi_{G0})$. Thus, both evolved populations of $\phi_{G0}$ and $\phi_{S0}$ are generalists with different emphasis on growth and survival.
@@
A few populations under growth-stress fall in the second quadrant, showing survival of $\phi_{S0}$ under stress conditions and extinction of $\phi_{G0}$. In other words, a strain evolved under stressful conditions survives under growth-stress cycles, while one evolved under growth cycles does not and is a specialist. However, as mentioned, the counterparts of these populations under solely growth cycles exist in the upper right quadrant. This shows that even though a population of $\phi_{S0}$ evolves to exist under stress conditions when $\phi_{G0}$ cannot, under solely growth cycles the two populations have comparable long term growth rates and may coexist. This further suggests that despite the growth-survival trade-off, populations evolved under growth-stress may not show significant trade-offs with their fitness under only growth cycles.
Some populations shift to the lower left quadrant in Figure 6 under growth-stress cycles, implying neither $\phi_{G0}$ or $\phi_{S0}$ are able to grow under stress. This occurs at lower values of $g_{0}$, where the low fecundity does not allow the population to recover its numbers after the significant decrease in its population size upon exposure to stress. Increasing both $g_{0}$ and $T$ can help the population recover and reduce this effect.
We note that while here we are considering a single optimum phenotype, in real world situations we expect to see a population with a distribution in phenotypes. The width of this population distribution is set by the balance between the variance in fitness mutation effects and the effective selection strength. Thus, if the population width is large enough to accommodate different phenotypes and hence different long term growth rates, coexistence of two strains with similar positive long term growth rates may be possible. This could happen in two ways. 1) Under conditions where one of the strains is optimal, but the population distribution is broad enough to accommodate the second. This is made possible by an increase in the variance of mutational effects, $\sigma_{m}$. 2) Under conditions when both strains are suboptimal and hence have comparable fitness. Since both strains can be generalists as the populations in the first quadrant of Figure 8, this means that in cases where the emphasis on lag time and growth rate is different than the conditions of evolution, coexistence may be possible. The $x$ and $y$ axis of Figure 8 represent the fitness of each strain for an environment with the same growth time, $T$, at which the strains are evolved to maximize fitness.
We test the second possibility by simulating the competitive growth of the two optimal strains with a common finite resource. We calculate the optimal phenotypes $\phi_{G0}$ and $\phi_{S0}$ using the parameter values used to generate Figure 6, with $g_{0}=40$ and $T=22$. We further calculate their respective growth rates, survival probabilities and lag times using equations 1. 3 and 4. Let $N_{i}$ (where $i=(1,2)$) be the biomass of strain i at each simulation time step. Our simulations start with equal and small biomass values of two strains (taken to be $N_{i}=0.1$) in a finite resource, $r_{0}$. At the beginning of the simulation, both strains are assumed to be in lag phase. At each time step we test whether both populations continue to be in their lag phase. While a population is in lag phase, its biomass does not increase but the resource is depleted by a small amount proportional to the biomass. Once either strain crosses its calculated lag time we assume it follows the Monod equation for effective growth rate in a finite resource,
$g_{i}=g_{i}^{0}\frac{r}{K_{i}+r},$ (9)
where $i=(1,2)$, $g_{i}^{0}$ is the maximum growth rate of strain i calculated using Equation 1 and the constant $K_{i}$ is the value of resource amount $r$ at which the growth rate of strain i falls to half its maximum value. At each time step of the simulation we update the biomass of both strains by adding $g_{i}N_{i}$, as well as the resource which reduces by an amount $\delta r_{i}$ (where $i=(1,2)$) per unit biomass of each strain produced. We allow each strain to grow till the amount of resource per unit biomass of the strain falls below a predefined threshold ($r/N_{i}<10$), at which point the strain enters stationary phase. In stationary phase, growth is halted while the resource continues to be consumed at a smaller rate $0.1\delta r_{i}$, like in the lag phase. Note that the two strains enter both growth and stationary phases independently. Once the resource value drops below a second threshold ($r/N_{i}<5$), we put both strains either through stress by multiplying each strain’s biomass by its survival probability, or through dilution, by multiplying each strain’s biomass by $d=0.3$. This process is then repeated by replenishing the resource amount once again to $r_{0}$. We consider a strain extinct once its biomass falls to one-thousandth its initial value ($N_{i}<0.0001$).
The value of growth time $T$ in the evolution simulations in the previous section is set by the amount of initial resource added, $r_{0}$, although the exact relation between them is difficult to determine analytically. Hence it is not directly possible to know the value of $r_{0}$ for which our two strains are optimal. However, we do know that changing $r_{0}$ leads to a change upon the emphasis on growth and stress in the cycles.
Figures 7 (a) and (b) shows two biomass time series for two different values of $r_{0}$ for growth only cycles. We find that at resource values across a range, including $r_{0}=2.5$ and $10.0$, we find that coexistence between the two occurs, with the maximum biomass of the stress-type strain being lower than that of the growth-type strain. Here the stress-type strain is at a disadvantage because of a lower growth rate and larger lag time. Yet because of comparable positive long term growth rates (the upper right quadrant of Figure 6) the stress-type strain does not go extinct up to $40,000$ simulation time steps.
Figures 7 (c) and (d) shows similar time series plots for growth-stress cycles. At different resource values we are able to see (c) a case where only the stress-type strain survives, and (d) a case where both are able to coexist. Out-competition in (c) occurs because the long term growth rate of the growth-type strain is driven small or negative due to the inclusion of survival probability (the lower right quadrant of Figure 6). Upon increasing $r_{0}$ to $11$, we effectively increase the long term growth rate of both strainFigure III. It is interesting to note that the maximum biomass of the growth-type strain is still higher than that of the stress-type even though its long term growth rate is similar. This is because the growth rate of the growth-type strain is higher than that of the stress-type strain, while their long term growth rate remains similar. Thus, the amount of biomass both are able to produce in the first exponential phase, which is larger for the growth-type strain, remains maintained through subsequent cycles.
Figure 7: We test the possibility of coexistence of growth-type and stress-type strains in growth cycles (a-b) and growth-stress cycles (c-d). We first determine the optimal phenotype for both conditions using Equation 8 and its corresponding growth rate, lag time and survival probability. We then run resource limited growth competition simulations with two strains of the derived characteristics. (a) When the simulation has solely growth cycles, we see both strains survive for most resource values, $r_{0}=2.5$ is shown here. The growth-type strain has a higher biomass. (b) A higher value of $r_{0}=10.0$ with growth only cycles is shown here. (c) Starting with $r_{0}=8.0$ in growth-stress cycles, we find that the growth type strain goes extinct, while the stress type strain survives. (d) Starting with $r_{0}=11.0$ and growth-stress cycles, we find that that the two strains are able to coexist. For these results we use $l_{0}=30$, $\tau_{l}=1000$, $r_{0}=10^{5}$, $g_{0}=30$, $\tau_{g}=1200$, $\tau_{g}=1200$, $\gamma_{q}=100$, $\tau_{s}=850$ which gives us $\phi_{S0}=465.81$, $\phi_{G0}=390.74$, $g_{G0}=13.6437$, $g_{S0}=11.2300$, $t_{l,G0}=18.28$, $t_{l,S0}=16.03$, $s_{G0}=0.0001$ and $s_{S0}=0.0005$. The parameters for the coexistence simulation are $K=3$, $\delta r_{G0}=0.29$, $\delta r_{S0}=0.4$, the threshold for $r/N_{i}$ for transition into stationary phase is $10$. Both strains are started from $N_{i}=0.1$ and the extinction threshold is $N_{i}=0.0001$.
IV Discussion
@@
Growth-survival trade-offs, observed across microbial species, help explain key ecological phenomena like co-existence, population heterogeneity, and the preference for resource-scarce environments [36, 37]. We present a model of adaptation to extreme stress under such a trade-off. To our knowledge, this is a novel effort to model the role of quiescence in evolutionary adaptation to extreme stress. This approach is supported by evidence that stress response in yeast may follow a general strategy rather than evolving to specific stress conditions [1, 20, 38]. Studies in bacteria suggest that exposure to lower concentrations of antibiotics may prepare cells to take on the challenge of lethal concentrations [39, 40]. Further, the existence of the growth-lag trade-off suggests that starvation in stationary phase may be perceived a mild stress [41]. These observations support our hypothesis that quiescence, a response to the relatively mild stress of starvation, may provide a helpful step towards resistance to harsher stress.
@@
Previous work has demonstrated that yeast populations exposed to repeated cycles of growth-freeze-thaw increase their survival from $2\%$ to $70\%$ over the course of $100$ generations [9]. The evolved strains show key characteristics of quiescent cells including smaller, denser cells, increased cell wall thickness. Our model captures this evolutionary trajectory qualitatively at the phenotypic level. It provides a framework to explore its dependence on the frequency and strength of environmental change, as well as on features of the population such as its maximum growth rate and the nature of the variation of lag time with the population phenotype.
@@
We draw mutational fitness effects from a long-tailed Pearson distribution. The model is agnostic to the source of such mutations, which could be genetic or epigenetic in nature. Further, our framework does not speculate on the mechanism through which the memory of stress, or of intracellular trehalose produced in stationary phase can be inherited across generations. Interestingly, the whole genome sequencing of the evolved yeast strains in [9] does not show any mutations conserved across replicate lines. This is also supported by previous work [47]. Yet, a convergent phenotype emerges within the short period of $25$ generations, suggesting an epigenetic stress response. This response is still inheritable, as evidenced by the evolved phenotype persisting even in the absence of continued stress. It is possible that if quiescence, a generalized stress response mechanism, facilitates the adaptation to freeze thaw, that the adaptation mechanism is one that involves genetic assimilation [48, 49].
We assume a stochastic switch to quiescence, with the probability of quiescence increasing with increasing intracellular trehalose. In this way our system is similar to stochastic switching to a persistent state in bacterial populations, although for us the probability of switching is biased by trehalose concentration in stationary phase [50, 51, 52]. However, our model is distinct from most models of bacterial persistors since we postulate that our phenotype is inheritable and over time lends higher survival fractions. There is some support for the idea that that persistence frequency in bacteria is also an evolvable trait that adapts to both the duration and frequency of the antibiotic application[61, 54]. We assume that the subset of quiescent cells is then subject to probabilistic survival upon exposure to stress. This is a reasonable assumption given previous evidence that quiescence leads to higher stress tolerance [55]. Yet this assumption remains experimentally unconfirmed, and it is not certain whether the same quiescent subset of the population in stationary phase is selected from by subsequent stress.
@@
We calculate the long term growth rate of a population adapting to growth-stress cycles and use it to calculate the optimum phenotype for different evolution programs. Our analysis reveals that the duration and nature of the growth process between successive stress exposures directly affects the optimum phenotype, final survival, as well as the long term growth rate. We compare the fitness of populations that are optimum under solely growth cycles and under growth-stress cycles. We calculate the long term growth rate of these ’growth-type’ and ’stress-type’ populations both with and without stress. Depending on the evolution program, the populations can evolve to be specialists or generalists. In particular, the long term growth rates of some stress-type populations remain positive and comparable to the long term growth rates of growth-type populations, both with and without stress. Since the long term growth rate of both strains is positive and comparable, this suggests the possibility of coexistence. We test this possibility using simulations of competition between the two strains under a similar growth paradigm with different growth durations, and indeed find coexistence under certain conditions. This indicates that the overall fitness of the two strains does not show a strong trade-off despite the underlying growth-survival trade-off. Our simulations treat populations growing in a fixed finite resource, and this makes a direct comparison with the long term growth rate calculation, which relies on fixed finite time growth, difficult.
For a smaller range of lower values of $g_{0}$ and time of growth, $T$, we find that stress-type populations have positive long term growth rates in both types of environments, while growth-type populations do not survive cycles with stress exposure. Thus, for some evolution paradigms, populations do not show a significant decrease in their fitness in growth-only cycles while still increasing their fitness in cycles with stress exposure. This is because the overall fitness and long term growth rate depend not only on the relative strengths of the growth and survival functions but also on other characteristics of growth such as time, $T$ and maximum growth rate $g_{0}$. Several experimental studies have shown that fitness trade-offs are not ubiquitous in microbial systems [4, 56]. Our results demonstrate that the absence of more obvious fitness trade-offs does not preclude the existence of an underlying growth-survival trade-off.
@@
Our model is also reminiscent of recent work done on the population dynamics of bacteria in stochastic feast-famine cycles [58, 59]. Previous work which models growth and death processes both with growth-yield and growth-death trade offs shows that the ratio of the growth rate and death rate is the effective fitness of the population [58]. This is similar to our expression for the long term growth rate which combines survival probability with the growth rate. Further, they also find the duration of the feast and famine periods to be important to the eventual fate of the population. However, here we also take into account the added complexity of lag time along with growth and survival, all determined by a single phenotype. Other work has examined the optimal persistence strategy for bacteria in feast-famine cycles with the stochastic application of antibiotics [60] to find that ’triggered’ dormancy, such as what we have included in this model with quiescence, may be the optimal strategy when the stress of the antibiotic treatment is very severe.